# Fisher-Rao Gradient Flow: Geodesic Convexity and Functional Inequalities

### v1

@@ 4.3. An easier-to-verify sufficient condition
(4.21) $(x-1)f^{\prime}(x)\geq f(x)-f(1)=f(x),\qquad\forall x>0.$
With the above preparations, we prove that for $0<x<1<y$, Condition 4.2 holds with $\alpha=\alpha_{s}\delta$. This can be obtained from the following two inequalities.
@@
$\geq\alpha_{s}\bigl{(}\frac{1}{x}-1\bigr{)}+\delta f^{\prime\prime}(1)\bigl{(}1-\frac{1}{y}\bigr{)}$
$\geq\alpha_{s}\delta\bigl{(}\frac{1}{x}-\frac{1}{y}\bigr{)}\quad(\textrm{Using }f^{\prime\prime}(1)\geq\alpha_{s})$
[…]
The proof is complete. ∎
@@
Recall that in general, the gradient dominance condition can be implied by strong geodesic convexity of the energy functional. In Theorem 3.4, a sufficient condition for strong geodesic convexity is that $xf^{\prime}(x)$ is concave and $f^{\prime\prime}(1)>0$. If $xf^{\prime}(x)$ is concave, we can show that $x^{2}f^{\prime\prime}(x)\geq f^{\prime\prime}(1)$ for any $0<x\leq 1$. To do so, we rewrite $x^{2}f^{\prime\prime}(x)$ in terms of $xf^{\prime}(x)$ as $x^{2}f^{\prime\prime}(x)=x\bigl{(}xf^{\prime}(x)\bigr{)}^{\prime}-xf^{\prime}(x).$ If $xf^{\prime}(x)$ is concave, then $h(x):=\bigl{(}xf^{\prime}(x)\bigr{)}^{\prime}=xf^{\prime\prime}(x)+f^{\prime}(x)$ is nonincreasing. We have
@@
$\geq 0\quad(\textrm{Using }h(1)\leq h(x),h(\xi)).$
Therefore, Theorem 4.6 is satisfied with $\alpha_{s}=f^{\prime\prime}(1)$, so the gradient dominance holds. This establishes a connection between Theorem 4.6 and Theorem 3.4.
5. Functional Inequality: Dual Gradient Dominance Condition
The KL divergence is the most widely used energy functional for sampling; however, the negative results in Theorems 3.3 and 4.1 imply that the concepts of geodesic convexity and the gradient dominance condition are not sufficient for analyzing the convergence of the Fisher-Rao gradient flow of the KL divergence.
In this section, we introduce a dual gradient dominance condition, providing a more powerful functional inequality to understand the convergence behavior of the Fisher-Rao gradient flow associated with various $f$-divergences. To the best of our knowledge, such dual functional inequalities have not been explored in the Wasserstein setting, making this a novel and non-trivial discovery in the context of the Fisher-Rao geometry.
5.1. Non-increasing of general $f$-divergences
Take a general convex function ${\bar{f}}\in C^{2}(0,+\infty),{\bar{f}}(1)=0$, we consider the following $\bar{f}$-divergence
(5.1) $D_{\bar{f}}[\rho\|{\rho^{*}}]=\int{\rho^{*}}{\bar{f}}(\frac{\rho}{{\rho^{*}}})\mathrm{d}\theta.$
Along the Fisher-Rao gradient flow (2.3) with respect to $D_{f}$, we have that
(5.2) $\frac{\mathrm{d}}{\mathrm{d}t}D_{\bar{f}}[\rho_{t}\|{\rho^{*}}]=\int\partial_{t}\rho_{t}{\bar{f}}^{\prime}\bigl{(}\frac{\rho_{t}}{{\rho^{*}}}\bigr{)}\mathrm{d}\theta$
$=$ $\int\Bigl{(}-\rho_{t}f^{\prime}(\frac{\rho_{t}}{{\rho^{*}}})+\rho_{t}\mathbb{E}_{\rho_{t}}\bigl{[}f^{\prime}(\frac{\rho_{t}}{{\rho^{*}}})\bigr{]}\Bigr{)}{\bar{f}}^{\prime}\bigl{(}\frac{\rho_{t}}{{\rho^{*}}}\bigr{)}\mathrm{d}\theta$
@@ 5.1. Non-increasing of general $f$-divergences
$\leq$ $0$
This shows that any $D_{\bar{f}}$ is non-increasing along the Fisher-Rao gradient flow with respect to $D_{f}$. Here in the last inequality, we used the fact that both $f^{\prime}(x)$ and ${\bar{f}}^{\prime}(x)$ are non-decreasing functions.
5.2. Dual gradient dominance condition
Motivated by the non-increasing of $D_{\bar{f}}$ for any ${\bar{f}}$, in the following we introduce the dual gradient dominance condition. The idea is that we require the derivative in (5.2) to be lower bounded by a constant multiplying the sum of $D_{f}$ and $D_{\bar{f}}$, similar to the gradient dominance condition but applied to a different energy functional $D_{\bar{f}}$.
@@ 5.2. Dual gradient dominance condition
There exists $\alpha_{f}>0$ depending only on $f$, such that
(5.3) $\begin{split}-\int\Bigl{(}-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl{[}f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr{]}\Bigr{)}{\bar{f}}^{\prime}(\frac{\rho}{{\rho^{*}}})\mathrm{d}\theta\geq\alpha_{f}\Bigl{(}D_{\bar{f}}[\rho\|{\rho^{*}}]+D_{f}[\rho\|{\rho^{*}}]\Bigr{)}.\end{split}$
The above inequality holds for any positive densities $\rho$ and ${\rho^{*}}$, provided that $f^{\prime}(\frac{\rho}{{\rho^{*}}})$ is absolutely integrable with respect to $\rho$.
Remark 5.1.
If the dual gradient dominance condition holds along the gradient flow of $D_{f}$, we will get
@@
which implies the exponential convergence, i.e., $\mathcal{O}(\exp(-\alpha_{f}t))$, of $D_{f}[\rho_{t}\|{\rho^{*}}]+D_{\bar{f}}[\rho_{t}\|{\rho^{*}}]$. $\lozenge$
In the following, we will consider two special choices of ${\bar{f}}$:
  1. (1)
    ${\bar{f}}(x)=\frac{(x-1)^{2}}{x}$ (corresponding to the reverse $\chi^{2}$ divergence)
  2. (2)
    ${\bar{f}}(x)=xf(\frac{1}{x})$ (corresponding to the *-conjugate of $f$ [67]).
We prove the dual gradient dominance condition holds under mild assumptions on $f$. Specifically, the first choice leads to a general result, while the second choice leads to tighter results for specific $f$ including the KL divergence.
5.3. General results using the reverse $\chi^{2}$ divergence as the dual
We take ${\bar{f}}(x)=\frac{(x-1)^{2}}{x}$, the $f$-divergence associated with ${\bar{f}}$ is the reverse $\chi^{2}$ distance
@@ 5.3. General results using the reverse $\chi^{2}$ divergence as the dual
Along the Fisher-Rao gradient flow (2.3) with respect to $D_{f}$, the derivative (5.2) is given explicitly as
@@
For this choice of ${\bar{f}}=\frac{(x-1)^{2}}{x}$, the generalized gradient dominance condition holds, which follows from the following theorem.
@@
Given a convex function $f\in C^{2}(0,+\infty)$ with $f(1)=0$ and $f^{\prime\prime}(1)>0$, there exists an $\alpha_{f}>0$ that depends on $f$ only, such that
(5.4a) $-\int\Bigl{(}-\rho f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}+\rho\mathbb{E}_{\rho}\bigl{[}f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\bigr{]}\Bigr{)}\Bigl{(}1-\bigl{(}\frac{{\rho^{*}}}{\rho}\bigr{)}^{2}\Bigr{)}\mathrm{d}\theta\geq D_{f}[\rho\|{\rho^{*}}],$
(5.4b) $-\int\Bigl{(}-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl{[}f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr{]}\Bigr{)}\Bigl{(}1-\bigl{(}\frac{{\rho^{*}}}{\rho}\bigr{)}^{2}\Bigr{)}\mathrm{d}\theta\geq 2\alpha_{f}\chi^{2}[{\rho^{*}}\|\rho].$
@@
Remark 5.3.
As a consequence of Theorem 5.2, the Fisher-Rao gradient flow (2.3) associated with any $f$-divergence, assumed well-posed, converges exponentially fast such that
$D_{f}[\rho_{t}\|{\rho^{*}}]+\chi^{2}[{\rho^{*}}\|\rho_{t}]\leq e^{-\alpha_{f}t}\Bigl{(}D_{f}[\rho_{0}\|{\rho^{*}}]+\chi^{2}[{\rho^{*}}\|\rho_{0}]\Bigr{)}.$
The investigation of the well-posedness of the equation is left as a future work. $\lozenge$
Proof for Theorem 5.2.
We denote
$\mu^{2}=\int\frac{{\rho^{*}}^{2}}{\rho}\mathrm{d}\theta\geq 1,$
then, the left hand side of (5.4a) or (5.4b) can be rewritten as
(5.5) $\begin{split}-\int\Bigl{(}-\rho f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}&+\rho\mathbb{E}_{\rho}\bigl{[}f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\bigr{]}\Bigr{)}\Bigl{(}1-\frac{{\rho^{*}}^{2}}{\rho^{2}}\Bigr{)}\mathrm{d}\theta\\ =&-\int f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\frac{{\rho^{*}}^{2}}{\rho}\mathrm{d}\theta+\int\frac{{\rho^{*}}^{2}}{\rho}\mathrm{d}\theta\cdot\int{\rho}f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\mathrm{d}\theta\\ =&-\int f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\Bigl{(}\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\Bigr{)}\rho\mathrm{d}\theta\\ =&\int\Bigl{(}f^{\prime}\Bigl{(}\frac{1}{\mu}\Bigr{)}-f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\Bigr{)}\Bigl{(}\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\Bigr{)}\rho\mathrm{d}\theta.\end{split}$
To proceed, we rely in the following local inequality. Its proof is in Appendix B.
@@
Denote $\mu^{2}=\int\frac{{\rho^{*}}^{2}}{\rho}\mathrm{d}\theta$, $\mu>0$, then for convex $f$ and $\frac{\rho^{*}}{\rho}\in(0,+\infty)$, it holds that
(5.6) $\Bigl{(}f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{\rho}{{\rho^{*}}})\Bigr{)}\Bigl{(}\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\Bigr{)}\geq\mu(\frac{{\rho^{*}}}{\rho}-\mu)f^{\prime}(\frac{1}{\mu})-\frac{\mu^{2}{\rho^{*}}}{\rho}(f(\frac{1}{\mu})-f(\frac{\rho}{{\rho^{*}}})).$
Multiplying the local inequality (5.6) by $\rho$ and integrating over $\mathbb{R}^{d}$ leads to
@@
(5.7) $\geq$ $\int\Bigl{(}\mu({\rho^{*}}-\mu\rho)f^{\prime}(\frac{1}{\mu})-\mu^{2}{\rho^{*}}(f(\frac{1}{\mu})-f(\frac{\rho}{{\rho^{*}}}))\Bigr{)}\mathrm{d}\theta$
$=$ $\mu(1-\mu)f^{\prime}(\frac{1}{\mu})-\mu^{2}f(\frac{1}{\mu})+\mu^{2}\int{\rho^{*}}f\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\mathrm{d}\theta$
$\geq$ $D_{f}[\rho\|{\rho^{*}}].$
In the last inequality, we used the fact that $0=f(1)\geq f(\frac{1}{\mu})-(\frac{1}{\mu}-1)f^{\prime}(\frac{1}{\mu})$ by the convexity of $f$; note that $\mu\geq 1$. This completes the proof of (5.4a).
To show (5.4b), we first introduce another local inequality as follows; its proof is in Appendix B.
@@
(5.8) $\frac{f^{\prime\prime}(1)}{2}\leq f^{\prime\prime}(x),\quad\text{for any }x\in(\frac{1}{1+2\delta_{f}},\frac{1}{1-2\delta_{f}}).$
Moreover, there exists $\alpha_{f}^{\prime}>0$, such that
(5.9) $\bigl{(}f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{1}{x})\bigr{)}(x^{2}-\mu^{2})\geq\alpha_{f}^{\prime}(x-\mu)^{2}$
@@
We will employ the quantities $\delta_{f}$ and $\alpha^{\prime}_{f}$ defined in Lemma 5.5. We prove the dual gradient dominance condition (5.4b) for the following two scenarios, depending on whether $\mu$ is greater than $1+\delta_{f}$ or not.
@@
(5.10) $\bigl{(}f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr{)}(\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2})\geq\alpha_{f}^{\prime}(\frac{{\rho^{*}}}{\rho}-\mu)^{2}.$
Multiplying the local inequality (5.10) by $\rho$ and integrating over $\mathbb{R}^{d}$ leads to
(5.11) $\begin{split}&\int\Bigl{(}f^{\prime}\Bigl{(}\frac{1}{\mu}\Bigr{)}-f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\Bigr{)}\Bigl{(}\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\Bigr{)}\rho\mathrm{d}\theta\geq\alpha^{\prime}_{f}\int(\frac{{\rho^{*}}}{\rho}-\mu)^{2}\rho\mathrm{d}\theta\\ =&\alpha^{\prime}_{f}\int\frac{{\rho^{*}}^{2}}{\rho}+\rho\mu^{2}-2{\rho^{*}}\mu\mathrm{d}\theta=2\alpha^{\prime}_{f}\mu(\mu-1)\quad(\textrm{using }\mu\geq 1)\\ \geq&\alpha^{\prime}_{f}(\mu+1)(\mu-1)=\alpha^{\prime}_{f}\chi^{2}[{\rho^{*}}\|\rho].\\ \end{split}$
When $\mu\geq 1+\delta_{f}$, utilizing (5.7) yields
@@
(5.12) $\begin{split}&\int\Bigl{(}f^{\prime}\Bigl{(}\frac{1}{\mu}\Bigr{)}-f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\Bigr{)}\Bigl{(}\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\Bigr{)}\rho\mathrm{d}\theta\\ \geq&\mu(1-\mu)f^{\prime}(\frac{1}{\mu})-\mu^{2}f(\frac{1}{\mu})+\mu^{2}\int{\rho^{*}}f\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\mathrm{d}\theta\\ \geq&\mu(1-\mu)f^{\prime}(\frac{1}{\mu})-\mu^{2}f(\frac{1}{\mu}).\end{split}$
@@
where in the last inequality, we used that the fact that the $f$-divergence is non-negative. Let denote $h(x)=(x-1)f^{\prime}(x)-f(x)$. Then for $x\in(0,1]$, $h(x)$ is nonnegative and non-increasing, since $h^{\prime}(x)=(x-1)f^{\prime\prime}(x)\leq 0$ and $h(1)=0$. Therefore, the inequality (LABEL:eq:generalized-gradient-dominant-condition-2b0) can be further carried on to obtain

### v2

@@ 4.3 An easier-to-verify sufficient condition
(75) $(x-1)f^{\prime}(x)\geq f(x)-f(1)=f(x),\qquad\forall x>0.$
With these preparations, we prove that for $0<x<1<y$, 4.2 holds with $\alpha=\alpha_{s}\delta$. This can be derived by multiplying the following two inequalities:
$\phantom{{}={}}f^{\prime}(y)-f^{\prime}(x)=\int_{x}^{1}f^{\prime\prime}(\xi){\rm d}\xi+\int_{1}^{y}f^{\prime\prime}(\xi){\rm d}\xi$
@@
$\geq\alpha_{s}\bigl(\frac{1}{x}-1\bigr)+\delta f^{\prime\prime}(1)\bigl(1-\frac{1}{y}\bigr)\quad\!\!(\textrm{using }\min\{\frac{y-1}{2},\delta\}\geq\frac{\delta}{y}(y-1)\textrm{ due to }y>1>2\delta)$
$\geq\alpha_{s}\delta\bigl(\frac{1}{x}-\frac{1}{y}\bigr)\quad(\textrm{using }\alpha_{s}>\alpha_{s}\delta\,\textrm{ and }\,f^{\prime\prime}(1)\geq\alpha_{s})$
@@
The proof is complete.
@@
Recall that in general, the gradient dominance condition is implied by strong geodesic convexity of the energy functional. In Theorem 3.3, a sufficient condition for strong geodesic convexity is that $xf^{\prime}(x)$ is concave and $f^{\prime\prime}(1)>0$. If $xf^{\prime}(x)$ is concave, we can show that $x^{2}f^{\prime\prime}(x)\geq f^{\prime\prime}(1)$ for any $0<x\leq 1$. To do so, we rewrite $x^{2}f^{\prime\prime}(x)$ in terms of $xf^{\prime}(x)$ as $x^{2}f^{\prime\prime}(x)=x\bigl(xf^{\prime}(x)\bigr)^{\prime}-xf^{\prime}(x).$ If $xf^{\prime}(x)$ is concave, then $h(x):=\bigl(xf^{\prime}(x)\bigr)^{\prime}=xf^{\prime\prime}(x)+f^{\prime}(x)$ is nonincreasing. We have
@@
$\geq 0\quad(\textrm{using }h(1)\leq h(x),h(\xi)\,\textrm{in}\,(0,1]).$
This implies that the assumption in Proposition 4.8 is satisfied with $\alpha_{s}=f^{\prime\prime}(1)$, so the gradient dominance holds. The argument here thus establishes a connection between Proposition 4.8 and Theorem 3.3.
5 Functional Inequality: Dual Gradient Dominance Condition
The KL divergence is the most widely used energy functional for sampling. However, the negative results in Theorems 3.1 and 4.1 indicate that the concepts of geodesic convexity and gradient dominance condition are insufficient for analyzing the convergence of the Fisher-Rao gradient flow of the KL divergence.
In this section, we introduce a dual gradient dominance condition. This condition provides a more powerful functional inequality for understanding the convergence behavior of Fisher-Rao gradient flows associated with various $f$-divergences. To the best of our knowledge, such dual functional inequalities have not been explored in the Wasserstein setting, making this a novel discovery in the context of Fisher-Rao geometry.
5.1 Non-increasing of general $f$-divergences
Given a general convex function ${\bar{f}}\in C^{2}(0,+\infty),{\bar{f}}(1)=0$, we consider the following $\bar{f}$-divergence
(76) $D_{\bar{f}}[\rho\|{\rho^{*}}]=\int{\rho^{*}}{\bar{f}}(\frac{\rho}{{\rho^{*}}})\mathrm{d}\theta.$
Along the Fisher-Rao gradient flow (25) with respect to $D_{f}$, we have that
(77) $\frac{\mathrm{d}}{\mathrm{d}t}D_{\bar{f}}[\rho_{t}\|{\rho^{*}}]=\int\partial_{t}\rho_{t}{\bar{f}}^{\prime}\bigl(\frac{\rho_{t}}{{\rho^{*}}}\bigr)\mathrm{d}\theta$
@@ 5.1 Non-increasing of general $f$-divergences
$\leq$ $0.$
Thus, any $D_{\bar{f}}$ is non-increasing along the Fisher-Rao gradient flow with respect to $D_{f}$. Here in the last inequality, we used the fact that both $f^{\prime}(x)$ and ${\bar{f}}^{\prime}(x)$ are non-decreasing functions.
5.2 Dual gradient dominance condition
Motivated by the non-increasing of $D_{\bar{f}}$ for any $\bar{f}$}, in the following we introduce the dual gradient dominance condition. The idea is that we require the derivative in (77) to be lower bounded by a constant multiplying the sum of $D_{f}$ and $D_{\bar{f}}$, similar to the gradient dominance condition but applied to a different energy functional $D_{\bar{f}}$.
@@ 5.2 Dual gradient dominance condition
The function $f$ is said to satisfy the dual gradient dominance condition if there exists a constant $\alpha_{f}>0$, depending only on $f$, such that
(78) $\begin{split}-\int\Bigl(-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl[f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr]\Bigr){\bar{f}}^{\prime}(\frac{\rho}{{\rho^{*}}})\mathrm{d}\theta\geq\alpha_{f}\Bigl(D_{\bar{f}}[\rho\|{\rho^{*}}]+D_{f}[\rho\|{\rho^{*}}]\Bigr),\end{split}$
for any positive densities $\rho$ and $\rho^{*}$}, provided that $f^{\prime}(\frac{\rho}{{\rho^{*}}})$ is absolutely integrable with respect to $\rho$.
If the dual gradient dominance condition holds along the gradient flow of $D_{f}$, then we have
$\frac{{\rm d}}{{\rm d}t}D_{\bar{f}}[\rho_{t}\|{\rho^{*}}]\leq-\alpha_{f}D_{\bar{f}}[\rho_{t}\|{\rho^{*}}],$
@@
which imply exponential convergence, i.e., $\mathcal{O}(\exp(-\alpha_{f}t))$, of $D_{\bar{f}}[\rho_{t}\|{\rho^{*}}]$, provided the initial condition $\rho_{0}$ satisfies $D_{\bar{f}}[\rho_{0}\|{\rho^{*}}]<\infty$. Similar comments apply to $D_{f}[\rho_{t}\|{\rho^{*}}]+D_{\bar{f}}[\rho_{t}\|{\rho^{*}}]$ provided $D_{f}[\rho_{0}\|{\rho^{*}}]+D_{\bar{f}}[\rho_{0}\|{\rho^{*}}]<\infty$. We note that commonly used particle-based initializations in sampling often violate this condition, while this condition may be satisfied by employing Gaussian [huang2022efficient, chen2023sampling], Gaussian mixture [lin2019fast, chen2024efficient, che2025stable], or kernel density approximations [maurais2024sampling, zhu2024kernel]. In the following, we will consider two special choices of $\bar{f}$}:
  1. 1.
    ${\bar{f}}(x)=\frac{(x-1)^{2}}{x}$, for which $D_{\bar{f}}$ corresponds to the reverse $\chi^{2}$ divergence.
  2. 2.
    ${\bar{f}}(x)=xf(\frac{1}{x})$, which corresponds to the *-conjugate of $f$ [osterreicher2002csiszar].
For both choices, we will prove the dual gradient dominance condition holds under certain assumptions on $f$. Specifically, the first choice leads to a general result, while the second choice leads to tighter results for specific $f$ including the KL divergence.
5.3 General results using the reverse $\chi^{2}$ divergence as the dual
We take ${\bar{f}}(x)=\frac{(x-1)^{2}}{x}$, the $f$-divergence associated with $\bar{f}$} is the reverse $\chi^{2}$ divergence
@@ 5.3 General results using the reverse $\chi^{2}$ divergence as the dual
Along the Fisher-Rao gradient flow (25) with respect to $D_{f}$, the derivative (77) is given explicitly as
@@
We have the following theorem:
@@
Given a convex function $f\in C^{2}(0,+\infty)$ with $f(1)=0$ and $f^{\prime\prime}(1)>0$, there exists an $\alpha_{\chi^{2},f}>0$ that depends on $f$ only, such that
(79a) $-\int\Bigl(-\rho f^{\prime}\Bigl(\frac{\rho}{{\rho^{*}}}\Bigr)+\rho\mathbb{E}_{\rho}\bigl[f^{\prime}\Bigl(\frac{\rho}{{\rho^{*}}}\Bigr)\bigr]\Bigr)\Bigl(1-\bigl(\frac{{\rho^{*}}}{\rho}\bigr)^{2}\Bigr)\mathrm{d}\theta\geq D_{f}[\rho\|{\rho^{*}}],$
(79b) $-\int\Bigl(-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl[f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr]\Bigr)\Bigl(1-\bigl(\frac{{\rho^{*}}}{\rho}\bigr)^{2}\Bigr)\mathrm{d}\theta\geq\alpha_{\chi^{2},f}\chi^{2}[{\rho^{*}}\|\rho].$
@@
a consequence of Theorem 5.1, the Fisher-Rao gradient flow (25) associated with any $f$-divergence will satisfy the dual gradient dominance condition with rate $\alpha_{f}=\frac{\min\{1,\alpha_{\chi^{2},f}\}}{2}$. Thus, assuming the flow is well-posed and using the discussions in LABEL:rmk:dual-gdc, we have that the flow converges exponentially fast:
$D_{f}[\rho_{t}\|{\rho^{*}}]+\chi^{2}[{\rho^{*}}\|\rho_{t}]\leq e^{-\frac{\min\{1,\alpha_{\chi^{2},f}\}}{2}t}\Bigl(D_{f}[\rho_{0}\|{\rho^{*}}]+\chi^{2}[{\rho^{*}}\|\rho_{0}]\Bigr).$
We require that the initial condition $\rho_{0}$ satisfies $D_{f}[\rho_{0}\|{\rho^{*}}]+\chi^{2}[{\rho^{*}}\|\rho_{0}]<\infty$.
Proof 5.2 (Proof for Theorem 5.1).
We denote $\mu^{2}=\int\frac{{\rho^{*}}^{2}}{\rho}\mathrm{d}\theta\geq 1,$ then, the left hand side of (79a) or (79b) can be rewritten as
(80) $\begin{split}-\int\Bigl(-\rho f^{\prime}\Bigl(\frac{\rho}{{\rho^{*}}}\Bigr)&+\rho\mathbb{E}_{\rho}\bigl[f^{\prime}\Bigl(\frac{\rho}{{\rho^{*}}}\Bigr)\bigr]\Bigr)\Bigl(1-\frac{{\rho^{*}}^{2}}{\rho^{2}}\Bigr)\mathrm{d}\theta\\ =&-\int f^{\prime}\Bigl(\frac{\rho}{{\rho^{*}}}\Bigr)\frac{{\rho^{*}}^{2}}{\rho}\mathrm{d}\theta+\int\frac{{\rho^{*}}^{2}}{\rho}\mathrm{d}\theta\int{\rho}f^{\prime}\Bigl(\frac{\rho}{{\rho^{*}}}\Bigr)\mathrm{d}\theta\\ =&-\int f^{\prime}\Bigl(\frac{\rho}{{\rho^{*}}}\Bigr)\Bigl(\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\Bigr)\rho\mathrm{d}\theta\\ =&\int\Bigl(f^{\prime}\Bigl(\frac{1}{\mu}\Bigr)-f^{\prime}\Bigl(\frac{\rho}{{\rho^{*}}}\Bigr)\Bigr)\Bigl(\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\Bigr)\rho\mathrm{d}\theta.\end{split}$
To proceed, we use the following lemma of a local inequality; its proof can be found in Appendix B.
@@
Given smooth densities $\rho$ and $\rho^{*}$, define $\mu^{2}=\int\frac{{\rho^{*}}^{2}}{\rho}\mathrm{d}\theta$, $\mu>0$. Then for convex $f$ and $\frac{\rho^{*}}{\rho}\in(0,+\infty)$, it holds that
(81) $\Bigl(f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{\rho}{{\rho^{*}}})\Bigr)\Bigl(\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\Bigr)\geq\mu(\frac{{\rho^{*}}}{\rho}-\mu)f^{\prime}(\frac{1}{\mu})-\frac{\mu^{2}{\rho^{*}}}{\rho}(f(\frac{1}{\mu})-f(\frac{\rho}{{\rho^{*}}})).$
Multiplying the local inequality (81) by $\rho$ and integrating over $\mathbb{R}^{d}$ leads to
@@
(82) $\geq$ $\int\mu({\rho^{*}}-\mu\rho)f^{\prime}\bigl(\frac{1}{\mu}\bigr)-\mu^{2}{\rho^{*}}\Bigl(f(\frac{1}{\mu})-f(\frac{\rho}{{\rho^{*}}})\Bigr)\mathrm{d}\theta$
$=$ $\mu(1-\mu)f^{\prime}(\frac{1}{\mu})-\mu^{2}f(\frac{1}{\mu})+\mu^{2}\int{\rho^{*}}f\bigl(\frac{\rho}{{\rho^{*}}}\bigr)\mathrm{d}\theta\geq D_{f}[\rho\|{\rho^{*}}].$
In the last inequality, we used the fact that $0=f(1)\geq f(\frac{1}{\mu})-(\frac{1}{\mu}-1)f^{\prime}(\frac{1}{\mu})$ by the convexity of $f$ and that $\mu\geq 1$. This completes the proof of (79a).
To show (79b), we introduce another local inequality with its proof is in Appendix B:
@@
(83) $\frac{f^{\prime\prime}(1)}{2}\leq f^{\prime\prime}(x),\quad\text{for any }x\in(\frac{1}{1+2\delta_{f}},\frac{1}{1-2\delta_{f}}).$
Moreover, there exists $\alpha_{\chi^{2},f}^{\prime}>0$, such that
(84) $\bigl(f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{1}{x})\bigr)(x^{2}-\mu^{2})\geq\alpha_{\chi^{2},f}^{\prime}(x-\mu)^{2}$
@@
We prove the dual gradient dominance condition (79b) separately for two different scenarios, depending on whether $\mu$ is greater than $1+\delta_{f}$ or not.
@@
(85) $\bigl(f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr)\bigl(\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\bigr)\geq\alpha_{\chi^{2},f}^{\prime}\bigl(\frac{{\rho^{*}}}{\rho}-\mu\bigr)^{2}.$
Multiplying the local inequality (85) by $\rho$ and integrating over $\mathbb{R}^{d}$ leads to
[…]
When $\mu\geq 1+\delta_{f}$, utilizing (82) yields
@@
(87) $\begin{split}&\int\Bigl(f^{\prime}\bigl(\frac{1}{\mu}\bigr)-f^{\prime}\bigl(\frac{\rho}{{\rho^{*}}}\bigr)\Bigr)\bigl(\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\bigr)\rho\mathrm{d}\theta\\ \geq&\mu(1-\mu)f^{\prime}(\frac{1}{\mu})-\mu^{2}f(\frac{1}{\mu})+\mu^{2}\int{\rho^{*}}f\Bigl(\frac{\rho}{{\rho^{*}}}\Bigr)\mathrm{d}\theta\\ \geq&\mu(1-\mu)f^{\prime}(\frac{1}{\mu})-\mu^{2}f(\frac{1}{\mu}).\end{split}$
@@
where in the last inequality, we used the fact that the $f$-divergence is non-negative. Let us denote $h(x)=(x-1)f^{\prime}(x)-f(x)$. Then for $x\in(0,1]$, $h(x)$ is nonnegative and non-increasing, since $h^{\prime}(x)=(x-1)f^{\prime\prime}(x)\leq 0$ and $h(1)=0$. Therefore, the inequality (LABEL:eq:generalized-gradient-dominant-condition-2b0) can be further carried on to obtain